OpenStudy (anonymous):

Find inverse of F(x) = x+1/x-4

OpenStudy (anonymous):

in know about how to switch values then multiply fracton but i'm stuck at seperating x from y

OpenStudy (anonymous):

(4x+1)/x-1

OpenStudy (anonymous):

so far i have xy -4x = y+1

OpenStudy (anonymous):

okay im not quite sure what you are asking but normally you solve for lets say x in terms of y then plug the answer to x back into the equation to solve for y

OpenStudy (anonymous):

yes you get x= y+1/y-4 multiply bottom to cancel out then you multiply other side then you get xy -4x = y+1 then how to seperate xy

OpenStudy (anonymous):

I have no idea anymore lol. I am just as lost as you. Sorry the answer above is correct though.

OpenStudy (anonymous):

it's not correct answer on my assignment it says 4x+1/x-1

OpenStudy (anonymous):

i got it correct now... let me type for you

OpenStudy (anonymous):

4x+1/x-1 = (4x+1)/x-1 parentheses in this case just organize

OpenStudy (anonymous):

\[f(x)= \frac{ x+1 }{ x-4 } x \neq4\] \[y=\frac{ x+1 }{ x-4 }\] \[xy-4y=x+1\] \[xy-x=4y+1\] \[x(y-1)=4y+1\] \[x=\frac{ 4y+1 }{ y-1 }\] Therefore now the inverse is \[y=\frac{ 4x+1 }{ x-1 }\]

OpenStudy (anonymous):

thanks soo much

OpenStudy (anonymous):

YOu are most welcomed